Pre-Final Course Grade – Abruzzese, Emily

Lab/Discussion Average: 93.19 Exam Average: 82.8 (lowest score counts half as much as better two) Problem Set Average: 94.8 (counted only if to your benefit) Course Average: 86 (based only on work to date)

Pre-Final Course Grade – Belanger, Caitlin

Lab/Discussion Average: 93.19 Exam Average: 70.8 (lowest score counts half as much as better two) Problem Set Average: 89.07 (counted only if to your benefit) Course Average: 76.61 (based only on work to date)

Pre-Final Course Grade – Conroy, Kathlyn

Lab/Discussion Average: 88.5 Exam Average: 68 (lowest score counts half as much as better two) Problem Set Average: 65.68 (counted only if to your benefit) Course Average: 71.42 (based only on work to date)

Pre-Final Course Grade – Corrado, Alexandra

Lab/Discussion Average: 88.5 Exam Average: 71.2 (lowest score counts half as much as better two) Problem Set Average: 74.45 (counted only if to your benefit) Course Average: 74.14 (based only on work to date)

Pre-Final Course Grade – Ferketic, Mary

Lab/Discussion Average: 94.9 Exam Average: 87.2 (lowest score counts half as much as better two) Problem Set Average: 87.45 (counted only if to your benefit) Course Average: 88.48 (based only on work to date)

Pre-Final Course Grade – Fimognari, Allison

Lab/Discussion Average: 97.28 Exam Average: 94.6 (lowest score counts half as much as better two) Problem Set Average: 89.3 (counted only if to your benefit) Course Average: 95.05 (based only on work to date)

Pre-Final Course Grade – Fournier, Rachel

Lab/Discussion Average: 97.28 Exam Average: 85.8 (lowest score counts half as much as better two) Problem Set Average: 92.35 (counted only if to your benefit) Course Average: 88.38 (based only on work to date)

Pre-Final Course Grade – Gallant, Monique

Lab/Discussion Average: 96.57 Exam Average: 80.8 (lowest score counts half as much as better two) Problem Set Average: 85.83 (counted only if to your benefit) Course Average: 83.77 (based only on work to date)

Pre-Final Course Grade – Gilbert, Richard

Lab/Discussion Average: 96.57 Exam Average: 68 (lowest score counts half as much as better two) Problem Set Average: 85.73 (counted only if to your benefit) Course Average: 74.61 (based only on work to date)

Pre-Final Course Grade – Hage, Brandon

Lab/Discussion Average: 91.92 Exam Average: 95 (lowest score counts half as much as better two) Problem Set Average: 94.35 (counted only if to your benefit) Course Average: 94.47 (based only on work to date)

Pre-Final Course Grade – Hana, Brikena

Lab/Discussion Average: 91.92 Exam Average: 67.6 (lowest score counts half as much as better two) Problem Set Average: 87.08 (counted only if to your benefit) Course Average: 73.86 (based only on work to date)

Pre-Final Course Grade – Holm, Conner

Lab/Discussion Average: 77.79 Exam Average: 84.8 (lowest score counts half as much as better two) Problem Set Average: 80.43 (counted only if to your benefit) Course Average: 83.63 (based only on work to date)

Pre-Final Course Grade – John, Stephanie

Lab/Discussion Average: 61.26 Exam Average: 58.2 (lowest score counts half as much as better two) Problem Set Average: 82.03 (counted only if to your benefit) Course Average: 62.04 (based only on work to date)

Pre-Final Course Grade – Kico, Fioralda

Lab/Discussion Average: 61.26 Exam Average: 73.6 (lowest score counts half as much as better two) Problem Set Average: 77.88 (counted only if to your benefit) Course Average: 72.45 (based only on work to date)

Pre-Final Course Grade – Komorowski, Alison

Lab/Discussion Average: 94.51 Exam Average: 89.6 (lowest score counts half as much as better two) Problem Set Average: 92.43 (counted only if to your benefit) Course Average: 90.71 (based only on work to date)

Pre-Final Course Grade – Kopp, Christopher

Lab/Discussion Average: 95.23 Exam Average: 94.6 (lowest score counts half as much as better two) Problem Set Average: 96.88 (counted only if to your benefit) Course Average: 95.02 (based only on work to date)

Pre-Final Course Grade – Larson, Matthew

Lab/Discussion Average: 95.23 Exam Average: 81.8 (lowest score counts half as much as better two) Problem Set Average: 66.94 (counted only if to your benefit) Course Average: 84.04 (based only on work to date)

Pre-Final Course Grade – Maher, Alaina

Lab/Discussion Average: 91.76 Exam Average: 60.6 (lowest score counts half as much as better two) Problem Set Average: 79.07 (counted only if to your benefit) Course Average: 67.69 (based only on work to date)

Pre-Final Course Grade – Masury, Katherine

Lab/Discussion Average: 87.83 Exam Average: 80.2 (lowest score counts half as much as better two) Problem Set Average: 88.14 (counted only if to your benefit) Course Average: 82.42 (based only on work to date)

Pre-Final Course Grade – McLaughlin, Tori

Lab/Discussion Average: 90.57 Exam Average: 82.4 (lowest score counts half as much as better two) Problem Set Average: 80.8 (counted only if to your benefit) Course Average: 83.76 (based only on work to date)

Pre-Final Course Grade – Meegan, Erin

Lab/Discussion Average: 90.57 Exam Average: 78.8 (lowest score counts half as much as better two) Problem Set Average: 69.58 (counted only if to your benefit) Course Average: 80.76 (based only on work to date)

Pre-Final Course Grade – Nguyen, Julie-Jennifer

Lab/Discussion Average: 90.57 Exam Average: 76 (lowest score counts half as much as better two) Problem Set Average: 83.3 (counted only if to your benefit) Course Average: 79.12 (based only on work to date)

Pre-Final Course Grade – O’Brien, Kevin

Lab/Discussion Average: 91.38 Exam Average: 92.4 (lowest score counts half as much as better two) Problem Set Average: 84.19 (counted only if to your benefit) Course Average: 92.23 (based only on work to date)

Pre-Final Course Grade – O’Connell, Katie

Lab/Discussion Average: 91.38 Exam Average: 59.8 (lowest score counts half as much as better two) Problem Set Average: 87.1 (counted only if to your benefit) Course Average: 68.21 (based only on work to date)

Pre-Final Course Grade – Oram, Thomas

Lab/Discussion Average: 88.97 Exam Average: 82.8 (lowest score counts half as much as better two) Problem Set Average: 81.97 (counted only if to your benefit) Course Average: 83.83 (based only on work to date)

Pre-Final Course Grade – Quinn, Carter

Lab/Discussion Average: 89.76 Exam Average: 79.6 (lowest score counts half as much as better two) Problem Set Average: 46.96 (counted only if to your benefit) Course Average: 81.29 (based only on work to date)

Pre-Final Course Grade – Quinn, McLean

Lab/Discussion Average: 79.06 Exam Average: 76.2 (lowest score counts half as much as better two) Problem Set Average: 44 (counted only if to your benefit) Course Average: 76.68 (based only on work to date)

Pre-Final Course Grade – Read, Jennifer

Lab/Discussion Average: 95.73 Exam Average: 80 (lowest score counts half as much as better two) Problem Set Average: 76.97 (counted only if to your benefit) Course Average: 82.62 (based only on work to date)

Pre-Final Course Grade – Smith, Amanda

Lab/Discussion Average: 91.41 Exam Average: 92.2 (lowest score counts half as much as better two) Problem Set Average: 81.37 (counted only if to your benefit) Course Average: 92.07 (based only on work to date)

Pre-Final Course Grade – Teebagy, Patrick

Lab/Discussion Average: 91.41 Exam Average: 70.6 (lowest score counts half as much as better two) Problem Set Average: 64.83 (counted only if to your benefit) Course Average: 74.07 (based only on work to date)

Pre-Final Course Grade – Torres, Sasha

Lab/Discussion Average: 92.06 Exam Average: 73.8 (lowest score counts half as much as better two) Problem Set Average: 68.03 (counted only if to your benefit) Course Average: 76.84 (based only on work to date)

Pre-Final Course Grade – Tran, Thu

Lab/Discussion Average: 92.06 Exam Average: 76 (lowest score counts half as much as better two) Problem Set Average: 99.93 (counted only if to your benefit) Course Average: 81.71 (based only on work to date)

Pre-Final Course Grade – Turpin, Julianne

Lab/Discussion Average: 94.17 Exam Average: 86.8 (lowest score counts half as much as better two) Problem Set Average: 92.71 (counted only if to your benefit) Course Average: 88.7 (based only on work to date)

Pre-Final Course Grade – Westin, Miriam

Lab/Discussion Average: 94.17 Exam Average: 94.8 (lowest score counts half as much as better two) Problem Set Average: 91.6 (counted only if to your benefit) Course Average: 94.69 (based only on work to date)